OpenStudy (anonymous):

plzzzz help If the sin 30° is 1 over 2, then the cos ____° = _____. 60° ; 1 over 2 30° ; square root 2 over 2 60° ; square root 3 over 2 30° ; 1

ganeshie8 (ganeshie8):

hint : \(\large \sin(\theta) = \cos(90 - \theta) \)

OpenStudy (anonymous):

so would it be A?

ganeshie8 (ganeshie8):

yes A is the answer, but wait a sec, something is wrong..

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

nvm, it works for all angles, so we're good :)

OpenStudy (anonymous):

@ganeshie8 thank you sooooo much

ganeshie8 (ganeshie8):

np :)